# Prompting Techniques for Reducing Social Bias in LLMs through System 1 and System 2 Cognitive Processes

### v1

@@ 5 Results & Analysis
All of the standard prompting results alongside the best performing prompting technique results for each bias category and model combination are presented in Figure 3. Here we see that the Human Persona with the System 2 (HP System 2) prompting technique often yields the least stereotypical responses, but that is not universal across models and bias categories. It outperforms all other prompting techniques in 13 cases. Similarly, the Human Persona in conjunction with System 2 and CoT (HP System 2 + CoT) prompting outperforms other prompting techniques in 9 cases. Only in one case, for profession bias and the GPT-4 model, the standard prompt outperform any other prompting techniques.
@@
We also see from Figure 3 that the two open-weight models, Llama 2 and Mistral 7B, often have similar behaviors (with the exception of the religion category) and that the two more recent closed-weight models, GPT-4 and Gemini have similar behaviors. GPT-4 stands out in having the most cases where prompting variants make minimal improvement from the standard zero-shot prompt. This suggests that OpenAI has done some engineering on this front. Though it is unclear whether this is behind-the-scenes prompt modifications, analogous instruction finetuning of the model, or some other method to make the model robust to prompt variants. Next, we focus on each bias category.
@@
Ageism bias in Figure 3 illustrates that various prompting techniques improved the performance of all models, reducing stereotypical responses by approximately 2 to 5 percent across the board compared to standard prompting.
@@
From Figure 3, it is evident that using various prompting techniques improves the outcomes across all models, achieving up to 13 percent reduction in stereotypical responses for beauty bias. Notably, the Mistral 7B model exhibited the most promising results with the Human Persona System 2 (HP System 2) prompting technique. Furthermore, among the five models tested, four showed enhanced performance when prompted with HP System 2.
@@
In the results depicted in Figure 3, it is evident that employing various prompting techniques leads to improved outcomes across all models, resulting in up to a 9 percent reduction in stereotypical responses for beauty in profession bias. Particularly noteworthy is the performance of the Gemini model when utilizing System 2 prompting, which shows the most promising results in reducing beauty in profession bias.
@@
We also observe a gender bias improvement of up to 9 percent across various models. For instance, the HP System 2 + CoT prompting technique performs well with GPT-3.5. Additionally, for GPT-4 and Gemini, the HP System 2 prompting yields the best performance. Notably, with Mistral 7B, HP System 1 outperforms other techniques. So, System 1 with Human Persona also reduces bias.
@@
We observed a reduction in institutional bias across all models, although the percentage decrease was smaller compared to reductions in gender or beauty biases. Specifically, with Llama 2, we achieved around a 5 percent improvement when using HP System 1 for prompting.
@@
Regarding nationality bias, the overall pattern of reduction is consistent across all models, similar to other biases. The combination of HP System 2 with CoT delivers the best performance for GPT-3.5 and Mistral 7B. For GPT-4 and Llama 2, the best results are achieved using HP alongside CoT.
@@
We achieved up to a 10 percent reduction in bias for the profession category. All models except GPT-4 exhibited less bias compared to the results from standard prompts. However, for GPT-4, standard prompting performed the best. In the case of Llama 2 and Mistral 7B, the HP System 2 prompts were the most effective, whereas for GPT-3.5 and Gemini, the System 2 prompts yielded the best results.
@@
We observe a reduction in racial biases across all models, although the decrease is relatively small for GPT-4, akin to that observed with standard prompting techniques. In contrast, the Mistral 7B model using HP System 2 with CoT prompting shows a bias reduction of approximately 9 percent.
@@
We achieved a reduction in religious bias by up to 8 percent. Additionally, we observed reductions across all models, although the decrease in the Gemini model was relatively minor, nearly identical to that seen with standard prompting. For GPT-3.5, MP System 2 exhibited the best performance. For Llama 2, CoT prompting delivered the least stereotypical responses, marking the only scenario where CoT prompting outperformed other prompting techniques.
6 Which Cognitive System Does CoT Best Model?
@@
CoT Vs Standard 0.4755 0.0 Reject
CoT Vs System 1 0.4580 0.0 Reject
CoT Vs System 2 0.4339 0.0 Reject
HP CoT Vs HP System 1 0.4639 0.0 Reject
HP CoT Vs HP System 2 0.4423 0.0 Reject
MP CoT Vs MP System 1 0.4556 0.0 Reject
MP CoT Vs MP System 2 0.4372 0.0 Reject
@@
Now we further investigate whether CoT prompting is most similar to System 2 reasoning in LLMs. While Figure 2 shows that the stereotyping rate of CoT is most similar to System 1 prompts, these may be from different test items. Here we tackle this question directly by computing the Kendall $\tau$ coefficient (Kendall, 1938) between CoT-prompted responses and those of the other variants. We selected the Kendall $\tau$ ranked correlation because there is a natural order to anti-stereotypical, neutral, and stereotypical categorical values in our datasets. Table 2 lists these results. From this, we find that CoT prompting is most similar to the Standard zero-shot prompt, followed by System 1 prompting. In fact, it is most dissimilar to System 2 prompting. This pattern holds for the Human Persona and Machine Persona variants, where CoT is least correlated with the System 2 prompt variant.
This raises questions regarding how to interpret the effect that CoT prompting has on LLM behavior. While prior work has found that CoT prompting leads to better multi-step mathematical and formal reasoning capabilities (Wei et al., 2022; Yu et al., 2023; Wang et al., 2023), which aligns with System 2 in dual process theory, we do not see the same effect in social biases in this study.
@@ 8 Conclusion
Our study has contributed to the understanding and reduction of social biases in LLMs through prompting techniques grounded in dual process theory. By harnessing the cognitive frameworks of System 1 and System 2, as well as the incorporation of human-like personas, our research not only clarifies the role of these cognitive processes in LLMs but also demonstrates practical methods for reducing biases. Our findings reveal that System 2 prompts, particularly when combined with a Human Persona, consistently reduce stereotypical judgments across various social bias categories. This indicates a profound potential for combining analytical thought processes and personalized prompting to enhance the ethical performance of LLMs. Furthermore, our use of different models and bias datasets has allowed us to explore the diverse applications of these techniques, ensuring our results are robust and applicable across different contexts. These findings underscore the potential of sophisticated prompting strategies in enhancing the ethical aspects of AI, pointing towards a future where LLMs can assist in creating more inclusive digital environments. Through continued exploration and refinement of these methods, we anticipate further advancements in the responsible deployment of AI technologies.

### v3

@@ 5.2 Model- and Bias-specific Prompting Effects
We now turn to specific model-bias category combinations. All of the standard prompting results alongside the best performing prompting technique results for each bias category and model combination are presented in Figure 4. Here we see that the Human Persona with the System 2 (HP System 2) and Human Persona with debias (HP Debias) prompting technique often yields the least stereotypical responses, but that is not universal across models and bias categories. HP Debias outperforms all other prompting techniques in 14 cases. Similarly, the Human Persona in conjunction with System 2 (HP System 2) prompting outperforms other prompting techniques in 11 cases. Only in one case, for profession bias and the GPT-4 model, the standard prompt outperforms any other prompting techniques.
@@
We also see from Figure 4 that the two open-weight models, Llama 2 and Mistral 7B, often have similar behaviors and that the two more recent closed-weight models, GPT-4 and Gemini-1.0 have similar behaviors. GPT-4 stands out in having the most cases where prompting variants make minimal improvement from the standard zero-shot prompt. This suggests that OpenAI has done some engineering on this front. Though it is unclear whether this is behind-the-scenes prompt modifications, analogous instruction finetuning of the model, or some other method to make the model robust to prompt variants. Next, we focus on each bias category.
@@
We see no consistent prompt setting that performs best on ageism. Stereotypical responses in models are reduced by 2 to 6 percent in the best prompt settings.
@@
Prompt variants in our experiments show substantial improvements on beauty bias across all considered models—up to 19 percent reduction in stereotypical responses in Mistral-7B using the HP System 2 prompt. The remaining 4 models also show major improvements in beauty bias, all using the HP Debias prompt.
@@
Gemini-1.0 shows a 13 percent reduction in stereotypical responses for beauty in profession bias using System 2 Debias prompting. The best prompt setting is inconsistent for this bias category, with HP System 2 showing the largest bias reduction for GPT-4 and Llama 2 while HP Debias and HP System 2 Debias results in the largest bias reduction in GPT-3.5 and Mistral-7B, respectively.
@@
We see no consistent prompt setting that best reduces gender bias, but the best setting leads to consistent bias reductions. Interestingly, the open-weight models, Llama-2 and Mistral-7B show less stereotypical gender responses before explored prompt-based improvements than the other models after.
@@
Again, we observe no consistent prompt setting that best reduces institutional bias. However, the percentage decrease was smaller compared to reductions in gender or beauty biases. With Llama 2, we achieved about a 5 percent improvement when using HP System 1 for prompting.
@@
Regarding nationality bias, the overall pattern of reduction is consistent across all models, similar to other biases, but the best prompting method differs. The combination of HP System 2 with CoT delivers the best performance for GPT-3.5 and Mistral 7B. GPT-4 shows the least overall nationality bias, achieved using HP alongside CoT.
@@
We achieved up to an 11 percent reduction in bias for the profession category. Here we see the single bias-model combination where standard prompting performed best in GPT-4. The other models all show substantial improvements. GPT-3.5 and Gemini-1.0, System 2 prompts yielded the best results. For Mistral-7B, Hp System 2 and for Llama 2, HP Debias were the best.
@@
We observe a reduction in racial biases across all models, although the decrease is relatively small for GPT-4, akin to that observed with standard prompting techniques. In contrast, the Mistral 7B model using HP System 2 with CoT prompting shows a bias reduction of approximately 9 percent. The best-performing prompting technique for race is HP Debias for GPT-3.5 which reduces around 17 percent of stereotypical engagement. Here HP Debias was the best prompting technique in three of the five models.
@@
We achieved a reduction in religious bias by up to 13 percent. Additionally, we observed reductions across all models, although the decrease in the GPT-4 model was relatively minor. Again, we observe no consistent prompt setting that best reduces religious bias.
6 Does CoT Prompting Best Model System 2?
@@
CoT Vs Standard 0.476 0.0 Reject
CoT Vs System 1 0.458 0.0 Reject
CoT Vs System 2 0.434 0.0 Reject
HP CoT Vs HP System 1 0.464 0.0 Reject
HP CoT Vs HP System 2 0.442 0.0 Reject
MP CoT Vs MP System 1 0.456 0.0 Reject
MP CoT Vs MP System 2 0.437 0.0 Reject
@@
Now we further investigate whether CoT prompting is most similar to the way that LLMs model System 2 reasoning. While Figure 2 shows that the stereotyping rate of CoT is most similar to System 1 prompts, these may be from different test items. Here we tackle this question directly by computing the Kendall $\tau$ coefficient (Kendall, 1938) between CoT-prompted responses and those of the other variants. We use the Kendall $\tau$ ranked correlation because there is a natural order to anti-stereotypical, neutral, and stereotypical categorical values in our datasets. Table 3 lists these results. From this, we find that CoT prompting is most similar to the Standard zero-shot prompt, followed by System 1 prompting. In fact, it is most dissimilar to System 2 prompting. This pattern holds for the Human Persona and Machine Persona variants, where CoT is least correlated with the System 2 prompt variant.
Our study aligns with Nighojkar’s (2024) results showing that CoT does not specifically resemble System 2. Nighojkar (2024) found that CoT prompting leads to LLMs better modeling human behavior, whether that is System 1 or System 2 depending on which cognitive process the setting triggers. While prior work has found that CoT prompting leads to better multi-step mathematical and formal reasoning capabilities (Wei et al., 2022; Yu et al., 2023; Wang et al., 2023), that align with System 2 cognitive processes, the growing body of evidence suggests that this is because the formal reasoning setting contextualizes LLMs to generate text reflecting System 2 reasoning in people.
@@ 8 Conclusion
Our study has contributed to the understanding and reduction of social biases in LLMs through prompting techniques grounded in dual process theory. By harnessing the cognitive frameworks of System 1 and System 2, as well as the incorporation of human-like personas and debiasing prompts, our research not only clarifies the role of these cognitive processes in LLMs but also demonstrates practical methods for reducing biases. Our findings reveal that System 2 prompts, particularly when combined with a Human Persona, consistently reduce stereotypical judgments across various social bias categories. Biases were further reduced when using a debiasing prompt, which can be seen as a social bias-focused System 2 prompt, along with a human persona. This indicates a profound potential for combining analytical thought processes and personalized prompting to enhance the ethical performance of LLMs. Furthermore, our use of different models and bias datasets has allowed us to explore the diverse applications of these techniques, ensuring our results are robust and applicable across different contexts. These findings underscore the potential of sophisticated prompting strategies in enhancing the ethical aspects of AI, pointing towards a future where LLMs can assist in creating more inclusive digital environments. Through continued exploration and refinement of these methods, we anticipate further advancements in the responsible deployment of AI technologies.